# Human-Exoskeleton Kinematic Calibration to Improve Hand Tracking for Dexterous Teleoperation

### v2

@@ Abstract
Hand exoskeletons are critical tools for dexterous teleoperation and immersive manipulation interfaces, but achieving accurate hand tracking remains a challenge due to user-specific anatomical variability and donning inconsistencies. These issues lead to kinematic misalignments that degrade tracking performance and limit applicability in precision tasks. We propose a subject-specific calibration framework for exoskeleton-based hand tracking that estimates virtual link parameters through residual-weighted optimization. A data-driven approach is introduced to empirically tune cost function weights using motion capture ground truth, enabling accurate and consistent calibration across users. Implemented on the Maestro hand exoskeleton with seven healthy participants, the method achieved substantial reductions in joint and fingertip tracking errors across diverse hand geometries. Qualitative visualizations using a Unity-based virtual hand further demonstrate improved motion fidelity. The proposed framework generalizes to exoskeletons with closed-loop kinematics and minimal sensing, laying the foundation for high-fidelity teleoperation and robot learning applications.
@@ I Introduction
Achieving human-level dexterity in robotic manipulation remains a key challenge [1], as current systems still fall short in achieving the versatility needed for tool use and fine manipulation [2, 3].
Dexterous teleoperation, where humans control robotic hands to perform fine tasks, offers a practical pathway toward this goal, supporting applications from surgery and space exploration to robot learning from human demonstration [4]. Its success depends critically on accurate, low-latency hand tracking that reliably captures user motion.
Existing approaches fall into two main categories: vision-based systems and wearable devices [5, 6]. Vision-based methods such as optical motion capture or camera-based tracking [7, 8], provide non-invasive sensing but suffer from occlusions, lighting sensitivity, and workspace limits, and they lack haptic feedback for force-based interactions [9].
@@
In contrast, wearable devices, such as data gloves or exoskeletons, directly measure joint angles or fingertip motion [10, 11, 12], enabling operation in cluttered environments and providing force feedback. However, their accuracy declines due to user-specific anatomy, inconsistent donning, and slippage [13, 14] and resulting in discomfort in mechanical designs that directly align device joints with anatomical joints [15].
Self-alignment mechanisms [16, 17] mitigate kinematic incompatibility but often introduce non-linear correspondence between device and anatomy. To complement these mechanical solutions, recent research has focused on calibration strategies that explicitly compensate for anatomical variability, donning inconsistencies, and sensor misalignment [18].
Commercial devices still rely on manual, pose-based calibration routines that are labor-intensive, user-dependent, and can take up to 90 minutes per glove, limiting practicality. Heinrich et al. [19] highlighted the trade-off between calibration accuracy and complexity, motivating faster, more generalizable approaches. Recent work by Li et al. [20] further advanced shape- and subject-specific calibration through unified kinematic–anatomical models that enhance tracking accuracy and consistency across users.
Hybrid and sensor-fusion approaches have also been explored. Gosala et al. [21] combined visual and magnetic sensors, reducing joint-angle error to 11.4° but requiring both cameras and wearables. Kim et al. [22] improved fingertip accuracy via data-driven optimization, though it depended on subject-specific ground-truth data.
@@
Specifically, we implement and validate this approach on the Maestro hand exoskeleton (Fig. 1), which employs self-aligning closed-loop four-bar linkages for kinematic compatibility. While this architecture enhances comfort and wearability, it remains sensitive to differences in hand size, anatomy, slippage, and donning variability [14]. To address these effects, we develop a subject-specific, data-driven calibration framework that estimates virtual link parameters and residual weightings directly from user motion data, compensating for geometric and anatomical variability without requiring rigid mechanical alignment.
The calibration is formulated as an optimization problem that minimizes tracking error through a residual-weighted cost function combining joint and fingertip discrepancies. The data-driven weighting scheme allows adaptation to subject-specific biomechanics and sensor contributions, resulting in more accurate hand-pose estimation and improved tracking fidelity. We evaluate the framework on seven participants with diverse hand geometries and validate its performance both quantitatively and qualitatively using a Unity-based virtual hand visualization. Although demonstrated on the Maestro exoskeleton, the proposed framework establishes a general foundation for accurate, subject-specific calibration in sensorized hand exoskeletons used for teleoperation and learning-from-demonstration.
@@ II-A Hardware Platform
The Maestro hand exoskeleton, shown in Fig. 1, is used in this study as a representative case to present and validate our subject-specific calibration framework. Maestro enables motion tracking and haptic feedback for the thumb, index, and middle fingers via a cable-driven mechanism [23, 24]. The device comprises 16 joints instrumented with rotary potentiometers: five sensors per index and middle finger and six for the thumb, each including two redundant measurements.
@@ II-D Human-Exoskeleton Calibration Framework
The kinematic calibration procedure relies on optimizing a set of virtual link parameters by minimizing a weighted sum of joint angle errors with the solver fminuncfminunc from MATLAB. The performance of this optimization depends not only on the quality of the input data but also on how the error terms are weighted in the cost function. While the simulation-based sensitivity analysis, presented in II-C, offers a first-principles rationale for prioritizing proximal parameters, it does not capture real-world factors such as model inaccuracies and inter-subject variability. To address this, we performed a data-driven weight refinement using motion-capture validation (see Section III-A).
Although illustrated on the Maestro four-bar design, the same calibration principle applies to other hand-exoskeleton architectures, provided that the system includes a sufficient number of independent constraints, whether from redundant robot sensors, geometric closures or constrained hand reference poses, to identify the unknown kinematic parameters. These constraints guarantee that the optimization remains well-posed; further discussion on generalizability is provided in Section V.
@@ IV-A Quantitative Results
The optimal-weighted calibration consistently yielded the lowest MAE, followed by the even-weighted calibration, while the uncalibrated model exhibited the highest errors across all tested joints. On average, the optimal-weighted approach produced greater error reductions, particularly at the fingertip level, with the index finger showing the most consistent improvements. The thumb also showed improvement following calibration, albeit with greater variability, which can be attributed to anatomical variability and the complexity of its joint articulation. Occasional negative error reductions in the thumb and index MCP joints may reflect residual alignment errors or unmodeled mechanical coupling. Overall, the consistent error reductions and comparable optimal weight profiles across participants indicate that the proposed calibration method generalizes well to users with varying hand sizes.
@@ V Discussion
Previous studies have also examined calibration trade-offs and shape-specific adaptation in data-glove systems [19, 20], which align with our findings and further emphasize the need for efficient, subject-agnostic calibration strategies.
@@
In terms of performance, the proposed method achieves a mean absolute fingertip error of approximately 10 mm, comparable to or better than markerless vision-based systems that often exceed 10–20 mm [25]. Prior studies have reported tolerable fingertip errors up to 10-15 mm without perceptible loss of control fidelity [26, 12], supporting the feasibility of the proposed framework for teleoperation tasks.
While the calibration approach is demonstrated on the Maestro hand exoskeleton, the proposed calibration framework extends naturally to other wearable devices where tracking depends on human parameters and there are sufficient independent constraints. For instance, in this study, the optimization was performed on kinematic loops lacking redundant sensing, such as the thumb IP joint, where closed-loop geometric constraints were incorporated to provide sufficient constraints. The data-driven weight optimization generalizes across users, as similar weight distributions were observed among participants, allowing an averaged profile to be used. Motion-capture calibration may only be needed once per device to account for design-specific geometry and sensors. The only fundamental requirement is that the number of virtual link parameters to be optimized does not exceed the number of independent kinematic constraints derived from redundant sensing, geometric closure, or constrained reference poses.
Despite the promising results, this study presents several limitations. First, although the evaluation included continuous finger motion, the analysis primarily focused on short, isolated movement segments rather than long-duration or coordinated hand actions. While this enabled precise joint-level analysis, it did not fully capture the temporal complexity of real-world manipulation tasks. Second, we acknowledge that the participants represent hand sizes slightly below the population mean, highlighting the need for validation across a broader and more diverse range of users. Additionally, this evaluation focuses on relative error reduction rather than direct comparisons with other devices. Future work will include systematic benchmarking to establish unified accuracy standards for reliable hand-exoskeleton deployment in teleoperation. Future work will also explore online or self-adaptive calibration schemes, allowing real-time adjustment to changes in fit or sensor drift. Extending the framework to simultaneous multi-finger calibration could further enhance dexterity modeling and real-world usability across diverse robotic hands.
@@ VI Conclusion
In conclusion, this work presents a subject-specific calibration framework that improves the accuracy of exoskeleton-based hand tracking for teleoperation. Implemented on the Maestro hand exoskeleton, the framework significantly reduces tracking errors by embedding anatomical variability derived from representative users directly into the device’s kinematic model. By optimizing virtual link parameters and residual weights, it achieves consistently higher hand tracking fidelity. Although residual joint-angle errors of about 10° remain, the calibration enables precise tracking for fine manipulation and shared-autonomy tasks. The proposed method is readily generalizable to other hand-exoskeleton designs, supporting user-specific hand motion tracking.

### v3

@@ Abstract
Hand exoskeletons are critical tools for dexterous teleoperation and immersive manipulation interfaces, but achieving accurate hand tracking remains a challenge due to user-specific anatomical variability and donning inconsistencies. These issues lead to kinematic misalignments that degrade tracking performance and limit applicability in precision tasks. We propose a subject-specific calibration framework for exoskeleton-based hand tracking that estimates virtual link parameters through residual-weighted optimization. A data-driven approach is introduced to empirically tune cost function weights using motion capture ground truth, enabling accurate and consistent calibration across users. Implemented on the Maestro hand exoskeleton with seven healthy participants, the method achieved substantial reductions in joint and fingertip tracking errors across diverse hand geometries. Qualitative visualizations using a Unity-based virtual hand further demonstrate improved motion fidelity. While demonstrated on the Maestro exoskeleton, the framework may be extended in principle to other hand exoskeleton systems that provide comparable kinematic constraints and sensing capabilities.
@@ I Introduction
Achieving human-level dexterity in robotic manipulation remains a key challenge [zhang2025dexterous], as current systems still fall short in achieving the versatility needed for tool use and fine manipulation [yu2022dexterous, chen2023visual].
Dexterous teleoperation, where humans control robotic hands to perform fine tasks, offers a practical pathway toward this goal, supporting applications from surgery and space exploration to robot learning from human demonstration [si2021review]. Its success depends critically on accurate, low-latency hand tracking that reliably captures user motion.
Existing approaches fall into two main categories: vision-based systems and wearable devices [mizera2019evaluation, fu2022teleoperation]. Vision-based methods such as optical motion capture or camera-based tracking [vakunov2020mediapipe, mueller2017real], provide non-invasive sensing but suffer from occlusions, lighting sensitivity, and workspace limits, and they lack haptic feedback for force-based interactions [moon2023hand].
@@
In contrast, wearable devices, such as data gloves or exoskeletons, directly measure joint angles or fingertip motion [darvish2023teleoperation, leonardis2024hand], enabling operation in cluttered environments and providing force feedback. However, their accuracy declines due to user-specific anatomy, inconsistent donning, and slippage [xia2022hand, yousaf2023experimental] and resulting in discomfort in mechanical designs that directly align device joints with anatomical joints [schiele2009ergonomics]. Self-alignment mechanisms [wei2023adaptive, zhang2025doglove] mitigate kinematic incompatibility but often introduce non-linear correspondence between device and anatomy. To complement these mechanical solutions, recent research has focused on calibration strategies that explicitly compensate for anatomical variability, donning inconsistencies, and sensor misalignment [connolly2022improving].
Commercial devices still rely on manual, pose-based calibration routines that are labor-intensive, user-dependent, and can take up to 90 minutes per glove, limiting practicality. Heinrich et al. [heinrich2024comparison] highlighted the trade-off between calibration accuracy and complexity, motivating faster, more generalizable approaches. Recent work by Li et al. [li2025fsglove] further advanced shape- and subject-specific calibration through unified kinematic–anatomical models that enhance tracking accuracy and consistency across users.
Hybrid and sensor-fusion approaches have also been explored. Gosala et al. [gosala2021self] combined visual and magnetic sensors, reducing joint-angle error to 11.4° but requiring both cameras and wearables. Kim et al. [kim2023calibration] improved fingertip accuracy via data-driven optimization, though it depended on subject-specific ground-truth data.
@@
Specifically, we implement and validate this approach on the Maestro hand exoskeleton (Fig. 1), which employs self-aligning closed-loop four-bar linkages for kinematic compatibility. While this architecture enhances comfort and wearability, it remains sensitive to differences in hand size, anatomy, slippage, and donning variability [yousaf2023experimental]. To address these effects, we develop a subject-specific, data-driven calibration framework that estimates virtual link parameters and residual weightings directly from user motion data, compensating for geometric and anatomical variability without requiring rigid mechanical alignment.
The calibration is formulated as an optimization problem that minimizes tracking error through a residual-weighted cost function combining joint and fingertip discrepancies. The data-driven weighting scheme allows adaptation to subject-specific biomechanics and sensor contributions, resulting in more accurate hand-pose estimation and improved tracking fidelity. We evaluate the framework on seven participants with diverse hand geometries and validate its performance both quantitatively and qualitatively using a Unity-based virtual hand visualization. Although this work focuses on the Maestro exoskeleton, the framework provides a systematic approach for subject-specific calibration in sensorized hand exoskeletons, provided that similar kinematic constraints and sensing capabilities are available.
@@ II-A Hardware Platform
The Maestro hand exoskeleton, shown in Fig. 1, is used in this study as a representative case to present and validate our subject-specific calibration framework. Maestro enables motion tracking and haptic feedback for the thumb, index, and middle fingers via a cable-driven mechanism [agarwal2015index, agarwal2017design]. The device comprises 16 joints instrumented with rotary potentiometers: five sensors per index and middle finger and six for the thumb, each including two redundant measurements.
@@ II-D Human-Exoskeleton Calibration Framework
The kinematic calibration procedure relies on optimizing a set of virtual link parameters by minimizing a weighted sum of joint angle errors with the solver fminuncfminunc from MATLAB. The performance of this optimization depends not only on the quality of the input data but also on how the error terms are weighted in the cost function. While the simulation-based sensitivity analysis, presented in II-C, offers a first-principles rationale for prioritizing proximal parameters, it does not capture real-world factors such as model inaccuracies and inter-subject variability. To address this, we performed a data-driven weight refinement using motion-capture validation.
Although illustrated on the Maestro four-bar design, the calibration principle can be extended in principle to other hand–exoskeleton architectures that provide sufficient independent constraints from redundant sensing, loop closures, or reference poses. In practice, each device still requires its own cost-function formulation and parameter tuning. Further discussion on scope and limitations is provided in Section V.
@@ IV-A Quantitative Results
The optimal-weighted calibration consistently yielded the lowest MAE, followed by the even-weighted calibration, while the uncalibrated model exhibited the highest errors across all tested joints. On average, the optimal-weighted approach produced greater error reductions, particularly at the fingertip level, with the index finger showing the most consistent improvements. The thumb also showed improvement following calibration, albeit with greater variability, which can be attributed to anatomical variability and the complexity of its joint articulation. Occasional negative error reductions in the thumb and index MCP joints may reflect residual alignment errors or unmodeled mechanical coupling. Overall, the consistent error reductions across participants suggest that the calibration improved tracking across participants with different hand sizes.
@@ V Discussion
Previous studies have also examined calibration trade-offs and shape-specific adaptation in data-glove systems [heinrich2024comparison, li2025fsglove], which align with our findings and further emphasize the need for efficient, subject-agnostic calibration strategies.
@@
In terms of performance, the proposed method achieves a mean absolute fingertip error of approximately 10 mm, comparable to or better than markerless vision-based systems that often exceed 10–20 mm [abdlkarim2024methodological]. Prior studies have reported tolerable fingertip errors up to 10-15 mm without perceptible loss of control fidelity [handa2020dexpilot, leonardis2024hand], supporting the feasibility of the proposed framework for teleoperation tasks.
The data-driven weight search revealed similar trends across participants, suggesting that an averaged profile is a reasonable initialization, though device- and user-specific refinement remains necessary for accurate tracking.
While demonstrated on the Maestro hand exoskeleton, the calibration framework can be generalized to other wearable systems that provide comparable independent kinematic constraints. While in this study geometric loop closures enabled solving joints without redundant sensing (e.g., the thumb IP), the optimization remains well-posed as long as the number of identifiable parameters does not exceed the available independent constraints.
Despite the promising results, the calibrated model does not remove all discrepancies: residual joint-angle and fingertip errors persist due to model mismatch, sensing noise, and unmodeled compliance, and are most noticeable at the fingertips, where small joint deviations accumulate along the kinematic chain. A complementary fingertip-based or task-level calibration can therefore be used to further refine fingertip alignment for dexterous tasks that demand higher accuracy.
This study has several other limitations. First, although the evaluation included continuous finger motion, the analysis primarily focused on short, isolated movement segments rather than long-duration or coordinated multi-finger actions, limiting insight into more complex manipulation behaviors. Second, we acknowledge that the participants represent hand sizes slightly below the population mean, highlighting the need for validation across a broader and more diverse range of users. Third, the evaluation assessed relative error reduction rather than benchmarking against other devices or calibration methods; establishing unified accuracy standards will be an important direction for future work.
Future work includes online or self-adaptive calibration to account for changes in fit or sensor drift, as well as extensions to simultaneous multi-finger calibration to further improve dexterity modeling and real-world usability across diverse robotic hands.
@@ VI Conclusion
In conclusion, this work presents a subject-specific calibration framework that improves the accuracy of exoskeleton-based hand tracking for teleoperation. Implemented on the Maestro hand exoskeleton, the calibration significantly reduces tracking errors by optimizing virtual link parameters and residual weights. While residual joint-angle and fingertip discrepancies remain, the method improves motion fidelity and can support teleoperation tasks. By optimizing virtual link parameters and residual weights, it achieves consistently higher hand tracking fidelity. Although residual joint-angle errors of about 10° remain, the calibration enables precise tracking for fine manipulation and teleoperation tasks. Rather than being universally applicable, the approach provides a structured calibration procedure that can be adapted to other hand–exoskeleton designs with similar kinematic constraints.